### v1

# 12, Proskura st., Kharkov 61085,Ukraine

Abstract
We consider the linear vector Schrödinger equation subjected to quadratic constraints. We demonstrate that the resulting nonlinear system is closely related to the Ablowitz-Ladik hierarchy and use this fact to derive the $N$-soliton solutions for the discussed model.
1 Introduction.
In this paper we want to discuss some nonlinear and seemingly integrable model in which the nonlinearity arises from the imposed constraints. We follow the approach developed by Pohlmeyer who considered in [1] the linear wave equation under quadratic constraints. This approach, which has been generalized by various authors, leads to the so-called $\sigma$-models, which play an important role in modern mathematics and physics (see, e.g., [2, 3, 4, 5, 6, 7]).
[…]
$\mathcal{L}=i\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}-\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}+\lambda\left(\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}-1\right)$ (1.1)
where $\boldsymbol{\psi}$ is a two-dimensional complex vector which is a function of two real variables $t$ and $x$, $\boldsymbol{\psi}=\boldsymbol{\psi}(t,x)$, $\boldsymbol{\psi}^{\dagger}$ is its Hermitian conjugate and subscripts denote derivatives with respect to the corresponding variables. The Lagrange multiplier $\lambda(t,x)$ is introduced to met the constraint
$\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}=1.$ (1.2)
The subject of our study are the Euler–Lagrange equations for (1.1), which can be written as
$0=i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+\lambda\boldsymbol{\psi},\qquad\lambda=\mathop{\mbox{Im}}\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}+\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}.$ (1.3)
The key point of this work is to demonstrate that equations (1.3) can be ‘embedded’ into the Ablowitz-Ladik hierarchy (ALH). In section 2 we show how one can obtain solutions for (1.3) from solutions for the equations of the ALH. Such approach was used in, say, [8, 9, 10] and was shown to be rather useful when one wants to find some particular solutions because the ALH is one of the most well-studied integrable systems. In section 3 we derive the $N$-soliton solutions for our problem by modification of the already known solitons of the ALH.
2 ALH hierarchy.
The ALH was introduced in 1976 in [11] as an infinite number of differential equations, the most famous of which is the discrete nonlinear Schródinger equation.
Later, it has been reformulated as a system of a few functional equations generated by the Miwa shifts applied to the functions of an infinite number of arguments (see [12]). The Miwa shifts, denoted by $\mathbb{E}_{\xi}$, are defined as
$\mathbb{E}_{\xi}q(\boldsymbol{z})=q(\boldsymbol{z}+i[\xi])$ (2.1)
where
$q(\boldsymbol{z})=q(z_{1},z_{2},...)=q(z_{k})_{k=1,...,\infty}$ (2.2)
and
$q(\boldsymbol{z}+i[\xi])=q(z_{1}+i\xi,z_{2}+i\xi^{2}/2,...)=q(z_{k}+i\xi^{k}/k)_{k=1,...,\infty}.$ (2.3)
In these terms, the ALH can be formulated as the following set of equations:
$\begin{array}[]{l}0=\tau_{n}\mathbb{E}_{\xi}{\tau_{n}}-\tau_{n-1}\mathbb{E}_{\xi}{\tau_{n+1}}-\rho_{n}\mathbb{E}_{\xi}{\sigma_{n}},\\[5.69054pt] 0=\tau_{n}\mathbb{E}_{\xi}{\sigma_{n}}-\sigma_{n}\mathbb{E}_{\xi}{\tau_{n}}-\xi\tau_{n-1}\mathbb{E}_{\xi}{\sigma_{n+1}},\\[5.69054pt] 0=\rho_{n}\mathbb{E}_{\xi}{\tau_{n}}-\tau_{n}\mathbb{E}_{\xi}{\rho_{n}}-\xi\rho_{n-1}\mathbb{E}_{\xi}{\tau_{n}}.\end{array}\quad n\in\left(-\infty,...,\infty\right).$ (2.4)
Strictly speaking, the above equations constitute only a half of the hierarchy, which is known to consist of two similar sub-hierarchies (the so-called ‘positive’ and ‘negative’ flows). However, for our current purposes, we may restrict ourselves to (2.4).
Now, we will derive some consequences of (2.4), which we use below to solve our problem. Introducing, for a fixed value of $n$,
$n=0,$ (2.5)
four new functions,
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\frac{\sigma_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\frac{\rho_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=\frac{\tau_{1}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\frac{\tau_{-1}}{\tau_{0}},$ (2.6)
one can show that these functions satisfy
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\xi\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\mathbb{E}_{\xi}{u},\qquad\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=-\xi v\,\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}$ (2.7)
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=-\xi\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\mathbb{E}_{\xi}{u},\qquad\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\xi v\,\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}$ (2.8)
where
$u=\frac{\sigma_{1}}{\tau_{0}},\qquad v=\frac{\rho_{-1}}{\tau_{0}}$ (2.9)
together with the constraint
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1.$ (2.10)
Returning from the functional equations to the differential ones with variables $z_{1}$ and $z_{2}$ being replaced with $x$ and $t$,
$x=z_{1},\quad t=z_{2},$ (2.11)
one can show, by means of the expansion
$\mathbb{E}_{\xi}q=q+i\xi q_{x}+\frac{\xi^{2}}{2}\left(iq_{t}-q_{xx}\right)+O\left(\xi^{3}\right),$ (2.12)
that functions $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$, $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$, $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ satisfy
$i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}=u\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}},\qquad i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}=-v\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}},$ (2.13)
$i\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}=-u\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\qquad i\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}=v\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ (2.14)
and
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{t}=u_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}-u\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{t}=v_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}-v\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x},$ (2.15)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{t}=u\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-u_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{t}=v\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}-v_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}.$ (2.16)
Now, we introduce two 2-vectors,
$\boldsymbol{q}=(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}},\;\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}})^{\scriptscriptstyle T},\qquad\boldsymbol{r}=(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\;\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}})^{\scriptscriptstyle T}$ (2.17)
and rewrite the above equations in the vector form,
$\boldsymbol{q}_{x}=u\,\mathsf{\sigma}_{2}\,\boldsymbol{r},\qquad\boldsymbol{r}_{x}=-v\,\mathsf{\sigma}_{2}\,\boldsymbol{q}$ (2.18)
and
$\boldsymbol{q}_{t}=iuv\boldsymbol{q}+iu_{x}\,\mathsf{\sigma}_{2}\,\boldsymbol{r},\qquad\boldsymbol{r}_{t}=-iuv\,\boldsymbol{r}+iv_{x}\,\mathsf{\sigma}_{2}\,\boldsymbol{q},$ (2.19)
where $\sigma_{2}=$ $\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right)$.
The restriction (2.10) now becomes
$\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}=1.$ (2.20)
By a simple algebra one can obtain the following consequence of (2.18) and (2.19):
$i\boldsymbol{q}_{t}+\boldsymbol{q}_{xx}+\lambda\boldsymbol{q}=0,$ (2.21)
$-i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+\lambda\boldsymbol{r}^{\scriptscriptstyle T}=0$ (2.22)
with $\lambda=2uv$ as well as the identities
$\lambda=-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{xx}=i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{xx}\boldsymbol{q}$ (2.23)
[…]
$\mathcal{L}=i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}+\lambda\left(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{{q}}-1\right)$ (2.24)
describing the vector linear Schrödinger system under the bilinear restriction (2.20). In other words, we have demonstrated that starting from (2.4) one can obtain solutions of the ‘two-field’ version of our problem.
An interesting fact, which has no direct relevance to our problem, is that $u$ and $v$ defined in (2.9) satisfy
$\left\{\begin{array}[]{r}iu_{t}+u_{xx}+2u^{2}v=0,\\ -iv_{t}+v_{xx}+2uv^{2}=0.\end{array}\right.$ (2.25)
So, as a by-products, we have obtained solutions for the nonlinear Schrödinger equation (see section 9.2 of [12]).
Till now, the correspondence between the ALH and the model (2.24) was rather general: any solution for (2.4) provide solution for (2.21)–(2.23). However, not all of them can be used to obtain solutions with $\boldsymbol{q}$ and $\boldsymbol{r}$ being relates by, say, $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. One thus need some additional work. Moreover, we have to make some slightly nonstandard steps. The case is that the ‘natural’ involution for the ALH-like equations is $\rho_{n}=\pm\sigma_{n}^{*}$, $\tau_{n}=\tau_{n}^{*}$ where * stands for the complex conjugation. Clearly, such involution does not provide necessary relation between $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$. Nevertheless, this issue can be resolved. In the next section we construct $N$-soliton solutions for our problem by modifying the already known ones that have been derived earlier for the ALH and discuss the issue of involution in more detail.
3 $N$-soliton solutions.
The main part of the structure of the soliton solutions are the so-called ‘soliton matrices’, that satisfy the system of Sylvester equations
$\mathsf{L}\mathsf{A}-\mathsf{A}\mathsf{R}=|\alpha\rangle\langle a|,\qquad\mathsf{R}\mathsf{B}-\mathsf{B}\mathsf{L}=|\beta\rangle\langle b|$ (3.1)
and that have been repeatedly used in the framework of the Cauchy matrix approach (see, e.g., [13, 14, 15, 16]).
Here, $\mathsf{L}$ and $\mathsf{R}$ are constant diagonal complex matrices,
$\mathsf{L}=\mbox{diag}\left(L_{1},...,L_{N}\right),\qquad\mathsf{R}=\mbox{diag}\left(R_{1},...,R_{N}\right),$ (3.2)
$|\alpha\rangle$ and $|\beta\rangle$ are constant $N$-columns,
$|\alpha\rangle=\left(\alpha_{1},...,\alpha_{N}\right)^{\scriptscriptstyle T},\qquad|\beta\rangle=\left(\beta_{1},...,\beta_{N}\right)^{\scriptscriptstyle T},$ (3.3)
while $N$-rows $\langle a|$ and $\langle b|$
$\langle a|=\left(a_{1},...,a_{N}\right),\qquad\langle b|=\left(b_{1},...,b_{N}\right)$ (3.4)
depend on the coordinates and, in turn, determine the coordinate dependence of the $N\times N$ matrices $\mathsf{A}$ and $\mathsf{B}$.
Using the results of the paper [17], one can present functions (2.6) as
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle a|\mathsf{R}^{-1}\mathsf{F}|\beta\rangle,$ (3.5)
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\langle b|\mathsf{L}^{-1}\mathsf{G}|\alpha\rangle,$ (3.6)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|\alpha\rangle,$ (3.7)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1+\langle b|\mathsf{L}^{-1}\mathsf{G}\mathsf{A}|\beta\rangle$ (3.8)
where
$\mathsf{F}=\left(1+\mathsf{B}\mathsf{A}\right)^{-1},\qquad\mathsf{G}=\left(1+\mathsf{A}\mathsf{B}\right)^{-1}.$ (3.9)
It can be shown that, if the Miwa shifts are implemented as
$\mathbb{E}_{\xi}\langle a|=\langle a|\mathsf{J}_{\xi}^{-1},\qquad\mathbb{E}_{\xi}\langle b|=\langle b|\mathsf{K}_{\xi}$ (3.10)
with
$\mathsf{J}_{\xi}=1-\xi\mathsf{R}^{-1},\qquad\mathsf{K}_{\xi}=1-\xi\mathsf{L}^{-1}$ (3.11)
which clearly implies
$\mathbb{E}_{\xi}\mathsf{A}=\mathsf{A}\mathsf{J}_{\xi}^{-1},\qquad\mathbb{E}_{\xi}\mathsf{B}=\mathsf{B}\mathsf{K}_{\xi}$ (3.12)
then the functions (3.5)–(3.8) satisfy (2.7) and (2.8) (see Appendix). Moreover, these functions are solutions of all differential equations of the ALH (equations (2.13)–(2.16) included), provided the variables of the hierarchy, $z_{1},z_{2},...$, are introduced in accordance with the definition (3.10). Returning to our problem, this means that the $(x,t)$-dependency is governed by
$i\mathsf{A}_{x}=\mathsf{A}\mathsf{R}^{-1},\qquad i\mathsf{A}_{t}=\mathsf{A}\mathsf{R}^{-2}$ (3.13)
and
$i\mathsf{B}_{x}=-\mathsf{B}\mathsf{L}^{-1},\qquad i\mathsf{B}_{t}=-\mathsf{B}\mathsf{L}^{-2},$ (3.14)
with similar equations for $\langle a|$ and $\langle b|$. Summarizing, we can formulate the following result.
Proposition 3.1
Vectors $\boldsymbol{q}$ and $\boldsymbol{r}$ defined in (2.17) and (3.5)–(3.8) with
$\langle a(x,t)|=\langle a_{0}|\mathsf{E}_{A}(x,t),\quad\mathsf{A}(x,t)=\mathsf{A}_{0}\mathsf{E}_{A}(x,t),$ (3.15)
$\langle b(x,t)|=\langle b_{0}|\mathsf{E}_{B}(x,t),\quad\mathsf{B}(x,t)=\mathsf{B}_{0}\mathsf{E}_{B}(x,t),$ (3.16)
where $\langle a_{0}|=\left(a_{01},...,a_{0N}\right)$ and $\langle b_{0}|=\left(b_{01},...,b_{0N}\right)$ are arbitrary constant $N$-rows,
$\mathsf{A}_{0}=\Biggl{(}\frac{\alpha_{j}a_{0k}}{L_{j}-R_{k}}\Biggr{)}_{j,k=1,...,N},\qquad\mathsf{B}_{0}=\Biggl{(}\frac{\beta_{j}b_{0k}}{R_{j}-L_{k}}\Biggr{)}_{j,k=1,...,N}$ (3.17)
and
$\mathsf{E}_{A}(x,t)=\exp\left(-ix\mathsf{R}^{-1}-it\mathsf{R}^{-2}\right),$ (3.18)
$\mathsf{E}_{B}(x,t)=\exp\left(ix\mathsf{L}^{-1}+it\mathsf{L}^{-2}\right)$ (3.19)
satisfy equations (2.21)–(2.23) and constraint (2.20).
Thus, we have derived soliton solutions for the ‘two-field’ version of our problem, consisting of linear Schrödinger equations under the bilinear constraint (2.20).
The last step is to pass from solutions described in Proposition 3.1 to solution of our problem. We start with imposing some restrictions on the constants involved to ensure the relations
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}{\!}^{*},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}{\!}^{*}.$ (3.20)
It turns out that this can be achieved by taking
$R_{j}=L_{j}^{*},\quad\beta_{j}a_{0j}=(\alpha_{j}b_{0j})^{*},\qquad j=1,...,N.$ (3.21)
This, at first, implies $\mathsf{E}_{B}=\mathsf{E}_{A}^{*}$. Then, after rewriting (3.17) as $\mathsf{A}_{0}=\mathsf{D}_{\alpha}\mathsf{C}\mathsf{D}_{a}$ and $\mathsf{B}_{0}=\mathsf{D}_{\beta}\mathsf{C}^{*}\mathsf{D}_{b}$, where
$\mathsf{C}=\Biggl{(}\frac{1}{L_{j}-L_{k}^{*}}\Biggr{)}_{j,k=1,...,N}$ (3.22)
and $\mathsf{D}_{\alpha}$, $\mathsf{D}_{\beta}$, $\mathsf{D}_{a}$, $\mathsf{D}_{b}$ are diagonal matrices with elements $\alpha_{j}$, $\beta_{j}$, $a_{0j}$, $b_{0j}$ ($j=1,...,N$) correspondingly, one can reveal the following structure of the matrices $\mathsf{F}$ and $\mathsf{G}$:
$\mathsf{F}=\mathsf{D}_{\beta}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}\mathsf{D}_{\beta}^{-1},\qquad\mathsf{G}=\mathsf{D}_{\alpha}\left(\mathsf{1}+\mathsf{Y}\mathsf{Y}^{*}\right)^{-1}\mathsf{D}_{\alpha}^{-1},$ (3.23)
where
$\mathsf{Y}=\mathsf{C}\mathsf{E},\quad\mathsf{E}=\mathsf{E}_{A}\mathsf{D}_{\beta}\mathsf{D}_{a}.$ (3.24)
After some simple calculations, one can rewrite $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ as
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle 1|\mathsf{R}^{-1}\mathsf{\Omega}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle 1|(\mathsf{R}^{*})^{-1}\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle$ (3.25)
where
$\mathsf{\Omega}=\mathsf{E}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}$ (3.26)
and
$\langle 1|=\left(1,...,1\right),\quad|1\rangle=\left(1,...,1\right)^{T}.$ (3.27)
Looking at (3.25) one can note that $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ are solitons of different type: $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ is a so-called bright soliton, vanishing as $x\to\pm\infty$ (with $t$ being fixed), while $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ is a dark soliton ($\lim_{x\to\pm\infty}|\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}|=1$). However, this does not mean that this is true for any solution, because of the fact that both the equations and constraints are invariant under transformations $\boldsymbol{q}\to\mathsf{U}\boldsymbol{q}$, if $\mathsf{U}^{\dagger}\mathsf{U}=\mathsf{1}$. So, general $N$-soliton solution is a mixture of dark and bright solitons.
Now, we have all necessary to formulate the main result of this paper.
Proposition 3.2
Vectors $\boldsymbol{\psi}$ defined by
$\boldsymbol{\psi}=\mathsf{U}\left(\begin{array}[]{c}\langle\gamma^{*}|\mathsf{\Omega}|1\rangle,\\ 1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle\end{array}\right)$ (3.28)
where $\mathsf{U}$ is an arbitrary unitary matrix,
$\mathsf{\Omega}=\mathsf{E}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},\quad\mathsf{Y}=\mathsf{C}\mathsf{E},\quad\mathsf{E}=\exp\left(-i\mathsf{\Theta}\right),$ (3.29)
$\mathsf{C}$ defined in (3.22),
$\mathsf{\Theta}=\mathsf{\Theta}(x,t)=\mathrm{diag}\left(x(L^{*}_{j})^{-1}+t(L^{*}_{j})^{-2}+\delta_{j}\right)_{j=1,..,N}$ (3.30)
and $\langle\gamma|$ is the constant $N$-row,
$\langle\gamma|=\left(1/L_{1},...,1/L_{N}\right),$ (3.31)
solve the Euler–Lagrange equations (1.3). Elements of the matrix $\mathsf{U}$ together with the $3N$ constants $\mathrm{Re}\,L_{j}$, $\mathrm{Im}\,L_{j}$ and $\delta_{j}$ are the arbitrary parameters that determine the properties of the $N$-soliton solution.
4 Conclusion.
In this work we have established the relationship between the Schrödinger equation with the constraint and the ALH. As was mentioned in section 2, we considered only the ‘positive’ subhierarchy (2.4). As to the ‘negative’ subhierarchy, it can be shown that calculations similar to ones presented above lead to the set of solutions similar to the solutions described in the proposition 3.2.
A more interesting question is whether we can tackle with the approach of this paper the case on quadratic restrictions other than (1.2), for example ones given by
$\boldsymbol{\psi}^{\dagger}\sigma_{3}\boldsymbol{\psi}=1$ (4.1)
where $\sigma_{3}=\mbox{diag}\left(1,-1\right)$? The answer, which we present here without derivation, is ‘yes’. However, to do this one should start not with the bright solitons of the ALH, as in section 3, but with the dark ones [18, 19]. In some sense the signature of the matrix describing the applied constraints play the role of the sign in front of the nonlinear term in the nonlinear Schrödinger equation: it determines which kind solitons (bright or dark) exists in the system.
Finally we would like to add a short comment on the integrability of the system (1.3) which was not discussed in the paper. We cannot at present prove its integrability, for example, by developing the inverse scattering transform. However, we now know that (1) it is a reduction of equations of the integrable ALH and that (2) it possesses $N$-soliton solutions. These two facts are a strong indication that (1.3) is integrable. Nevertheless, we think that the work in this direction should be continued, and one of the first problems to solve is to find the conservation laws of the model, which may be a topic of the following study.
Appendix.
Here we demonstrate that ansatz (3.5)–(3.8) leads to the solutions of the (2.7) and (2.8)
As follows from (3.5) and (3.10),
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|b\rangle-\langle{a}|\mathsf{R}^{-1}\mathsf{F}|b\rangle$
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|b\rangle-\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}\mathsf{J}_{\xi}\mathsf{F}|b\rangle$
which can be rewritten as
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}|b\rangle$ (A.1)
where $\mathsf{X}$ is defined by
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}=\mathbb{E}_{\xi}{\mathsf{F}}-\mathsf{J}_{\xi}\mathsf{F}.$
Using the definition of $\mathsf{F}$ (3.9), (3.12) and then (3.1) one can obtain
$\mathsf{X}$ $=$ $1-\mathsf{J}_{\xi}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}$
$=$ $\xi\mathsf{R}^{-1}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}$
$=$ $\xi\mathsf{R}^{-1}+\xi\mathsf{B}\mathsf{L}^{-1}\mathsf{A}$
$=$ $\xi\mathsf{R}^{-1}\mathsf{F}^{-1}+\xi\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}.$
Substituting this expression into (A.1) and using (3.8) together with the identity $\mathsf{A}\mathsf{F}=\mathsf{G}\mathsf{A}$, one arrives at
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle+\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}\mathsf{F}|b\rangle$
$=$ $\xi\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\mathbb{E}_{\xi}{{u}}$
which is nothing but the first equation from (2.7).
In a similar way one can prove that $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ satisfies the first equation from (2.8).
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})|a\rangle-\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|a\rangle$ (A.2)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}|a\rangle$
where
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}=(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})-\mathsf{J}_{\xi}\mathsf{F}\mathsf{B}.$
Calculating $\mathsf{Y}$,
$\mathsf{Y}$ $=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}+(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}\mathsf{B}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}\mathsf{B}$
$=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}$
$=$ $\mathsf{B}\mathsf{K}_{\xi}-\mathsf{J}_{\xi}\mathsf{B}$
$=$ $\xi\mathsf{R}^{-1}\mathsf{B}-\xi\mathsf{B}\mathsf{L}^{-1}$
$=$ $-\xi\mathsf{R}^{-1}|b\rangle\langle b|\mathsf{L}^{-1},$
and substituting it in (A.2) one can obtain
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $-\xi\langle b|\mathsf{L}^{-1}\mathsf{G}|a\rangle\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|b\rangle$
$=$ $-\xi\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\mathbb{E}_{\xi}{{u}},$
which concludes the proof.
The rest of the equations (2.7) and (2.8) can be tackled in a similar way.