### v2

# 12, Proskura st., Kharkiv 61085,Ukraine

Abstract
We consider the linear vector Schrödinger equation subjected to quadratic constraints. We demonstrate that the resulting nonlinear system is closely related to the Ablowitz-Ladik hierarchy and use this fact to derive the $N$-soliton solutions for the discussed model. As an example of application of these results we present solitons of some vector nonlinear Schrödinger equation with gradient nonlinearity.
1 Introduction.
In this paper we want to discuss some nonlinear and seemingly integrable model in which the nonlinearity arises from the imposed constraints. We follow the approach developed by Pohlmeyer who considered in [1] the linear wave equation under quadratic constraints. This approach, which has been generalized by various authors, leads to the so-called $\sigma$-models, which play an important role in modern mathematics and physics (see, e.g., the books [2, 3]).
Here, we would like to find some solutions for the problem when similar constraints are applied to the linear Schrödinger equation. This problem is described by the action $\mathcal{S}=\int dx\,dt\,\mathcal{L}$ with the Lagrangian
$\mathcal{L}=i\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}-\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}+\lambda\left(\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}-1\right)$ (1.1)
where $\boldsymbol{\psi}$ is a two-dimensional complex vector which is a function of two real variables $t$ and $x$, $\boldsymbol{\psi}=\boldsymbol{\psi}(t,x)$, $\boldsymbol{\psi}^{\dagger}$ is its Hermitian conjugate and subscripts denote derivatives with respect to the corresponding variables. The Lagrange multiplier $\lambda(t,x)$ is introduced to meet the constraint
$\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}=1.$ (1.2)
The subject of our study are the Euler–Lagrange equations for (1.1), which can be written as
$\begin{array}[]{r}i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+\lambda\boldsymbol{\psi}=0,\\ -i\boldsymbol{\psi}^{\dagger}_{t}+\boldsymbol{\psi}^{\dagger}_{xx}+\lambda\boldsymbol{\psi}^{\dagger}=0\end{array}$ (1.3)
with
$\lambda=\mathop{\mbox{Im}}\boldsymbol{\psi}^{\dagger}\boldsymbol{\psi}_{t}+\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}.$ (1.4)
The key point of this work is to demonstrate that equations (1.3) and (1.4) can be ‘embedded’ into the Ablowitz-Ladik hierarchy (ALH). In section 2 we show how one can obtain solutions for (1.3) from solutions for the equations of the ALH. Such approach was used in, for example, [4, 5] and was shown to be rather useful when one wants to find some particular solutions because the ALH is one of the most well-studied integrable systems. In section 3 we derive the $N$-soliton solutions for our problem by modification of the already known solitons of the ALH. In section 4 we consider an example of possible applications of the obtained results and present solitons of some vector nonlinear Schrödinger equation (NLSE) with gradient nonlinearity.
2 Ablowitz-Ladik hierarchy.
The ALH was introduced in 1976 in [6] as an infinite number of differential equations, the most famous of which is the discrete nonlinear Schrödinger equation.
Later, it has been reformulated as a system of a few functional equations generated by the Miwa shifts applied to the functions of an infinite number of arguments. The Miwa shifts, denoted by $\mathbb{E}_{\xi}$, are defined as
$\mathbb{E}_{\xi}q(\boldsymbol{z})=q(\boldsymbol{z}+i[\xi])$ (2.1)
where
$q(\boldsymbol{z})=q(z_{1},z_{2},...)=q(z_{k})_{k=1,...,\infty}$ (2.2)
and
$q(\boldsymbol{z}+i[\xi])=q(z_{1}+i\xi,z_{2}+i\xi^{2}/2,...)=q(z_{k}+i\xi^{k}/k)_{k=1,...,\infty}.$ (2.3)
In these terms, the ALH can be formulated as the following set of equations:
[…]
Strictly speaking, the above equations constitute only a half of the hierarchy, which is known to consist of two similar sub-hierarchies (the so-called ‘positive’ and ‘negative’ flows). However, for our current purposes, we may restrict ourselves to (2.4).
Now, we will derive some consequences of (2.4), which we use below to solve our problem.
First we introduce, for a fixed value of $n$,
$n=0,$ (2.5)
four new functions,
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\frac{\sigma_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\frac{\rho_{0}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=\frac{\tau_{1}}{\tau_{0}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\frac{\tau_{-1}}{\tau_{0}}.$ (2.6)
The original Ablowitz-Ladik equations were formulated in terms of $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ (with the $n$-dependence restored), and the first two equations in (2.6) are the standard way to introduce the tau-functions in order to arrive at bilinear equations (2.4). The second pair of functions, $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$, usually was not considered (or even introduced) in the framework of the Ablowitz-Ladik equations. However, as we see in what follows, they are a ‘natural’ complement to $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ and will play an essential role in this work.
One can show that functions defined in (2.6) satisfy
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\xi\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\mathbb{E}_{\xi}{u},\qquad\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=-\xi v\,\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}$ (2.7)
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=-\xi\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\mathbb{E}_{\xi}{u},\qquad\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\xi v\,\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}$ (2.8)
where
$u=\frac{\sigma_{1}}{\tau_{0}},\qquad v=\frac{\rho_{-1}}{\tau_{0}}$ (2.9)
together with the constraint
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1.$ (2.10)
Returning from the functional equations to the differential ones with variables $z_{1}$ and $z_{2}$ being replaced with $x$ and $t$,
$x=z_{1},\quad t=z_{2},$ (2.11)
one can show, by means of the expansion
$\mathbb{E}_{\xi}q=q+i\xi q_{x}+\frac{\xi^{2}}{2}\left(iq_{t}-q_{xx}\right)+O\left(\xi^{3}\right),$ (2.12)
that functions $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$, $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$, $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ satisfy
$i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}=u\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}},\qquad i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}=-v\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}},$ (2.13)
$i\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}=-u\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\qquad i\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}=v\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ (2.14)
and
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{t}=u_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}-u\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{t}=v_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}-v\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x},$ (2.15)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{t}=u\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-u_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{t}=v\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}-v_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}.$ (2.16)
Now, we introduce two 2-vectors,
$\boldsymbol{q}=(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}},\;\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}})^{\scriptscriptstyle T},\qquad\boldsymbol{r}=(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}},\;\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}})^{\scriptscriptstyle T}$ (2.17)
and rewrite the above equations in the vector form,
$\boldsymbol{q}_{x}=u\,\mathsf{\sigma}_{2}\,\boldsymbol{r},\qquad\boldsymbol{r}_{x}=-v\,\mathsf{\sigma}_{2}\,\boldsymbol{q}$ (2.18)
and
$\boldsymbol{q}_{t}=iuv\boldsymbol{q}+iu_{x}\,\mathsf{\sigma}_{2}\,\boldsymbol{r},\qquad\boldsymbol{r}_{t}=-iuv\,\boldsymbol{r}+iv_{x}\,\mathsf{\sigma}_{2}\,\boldsymbol{q},$ (2.19)
where $\sigma_{2}=$ $\left(\begin{array}[]{cc}0&-i\\ i&0\end{array}\right)$.
The restriction (2.10) now becomes
$\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}=1.$ (2.20)
By a simple algebra one can obtain the following consequence of (2.18) and (2.19):
$i\boldsymbol{q}_{t}+\boldsymbol{q}_{xx}+\lambda\boldsymbol{q}=0,$ (2.21)
$i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+\lambda\boldsymbol{r}^{\scriptscriptstyle T}=0$ (2.22)
with
$\lambda=2uv$ (2.23)
as well as the identities
$\lambda=-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{xx}=i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{xx}\boldsymbol{q}.$ (2.24)
It is easy to see that equations (2.21) and (2.22) together with (2.24) are the Euler–Lagrange equations for the action $\mathcal{S}=\int dx\,dt\;\mathcal{L}$ with the Lagrangian
$\mathcal{L}=i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}+\lambda\left(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{{q}}-1\right)$ (2.25)
describing the vector linear Schrödinger system under the bilinear restriction (2.20). In other words, we have demonstrated that starting from (2.4) one can obtain solutions of the ‘two-field’ version of our problem.
It should be noted that problem with the Lagrangian (2.25) is more general than one with (1.1) and our original problem is a reduction of the former: $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. In physical applications, models involving $\boldsymbol{\psi}$ and $\boldsymbol{\psi}^{\dagger}$ appear more often than similar models with two independent fields, as $\boldsymbol{q}$ and $\boldsymbol{r}$ in our case. However it is a common practice even in the cases of problems with complex fields, like, for example, the NLSE, consider $\boldsymbol{\psi}$ and $\boldsymbol{\psi}^{\dagger}$ as distinct variables because most part (but surely not all) of the calculations depend on the algebraic structure of the equations and not on the involution $\boldsymbol{\psi}\leftrightarrow\boldsymbol{\psi}^{\dagger}$.
An interesting fact, which has no direct relevance to our problem, is that $u$ and $v$ defined in (2.9) satisfy
$\left\{\begin{array}[]{r}iu_{t}+u_{xx}+2u^{2}v=0,\\ -iv_{t}+v_{xx}+2uv^{2}=0.\end{array}\right$. (2.26)
(we give a proof of this statement in Appendix A). So, as a by-products, we have obtained solutions for the NLSE.
Till now, the correspondence between the ALH and the model (2.25) was rather general: any solution for (2.4) provides solution for (2.21)–(2.24). However, not all of them can be used to obtain solutions with $\boldsymbol{q}$ and $\boldsymbol{r}$ being related by $\boldsymbol{r}^{\scriptscriptstyle T}=\boldsymbol{q}^{\dagger}$. One thus needs some additional work. Moreover, we have to make some slightly nonstandard steps. The case is that the ‘natural’ involution for the ALH-like equations is $\rho_{n}=\pm\sigma_{n}^{*}$, $\tau_{n}=\tau_{n}^{*}$ where * stands for the complex conjugation. Clearly, such involution does not provide necessary relation between $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$. Nevertheless, this issue can be resolved. In the next section we construct $N$-soliton solutions for our problem by modifying the already known ones that have been derived earlier for the ALH and discuss the issue of involution in more detail.
3 $N$-soliton solutions.
The main part of the structure of the soliton solutions are the so-called ‘soliton matrices’, that satisfy the system of the Sylvester equations
$\mathsf{L}\mathsf{A}-\mathsf{A}\mathsf{R}=|\alpha\rangle\langle a|,\qquad\mathsf{R}\mathsf{B}-\mathsf{B}\mathsf{L}=|\beta\rangle\langle b|$ (3.1)
and that have been repeatedly used in the framework of the Cauchy matrix approach (see, e.g., chapter 9 of [7]).
Here, $\mathsf{L}$ and $\mathsf{R}$ are constant diagonal complex matrices,
$\mathsf{L}=\mbox{diag}\left(L_{1},...,L_{N}\right),\qquad\mathsf{R}=\mbox{diag}\left(R_{1},...,R_{N}\right),$ (3.2)
$|\alpha\rangle$ and $|\beta\rangle$ are constant $N$-columns,
$|\alpha\rangle=\left(\alpha_{1},...,\alpha_{N}\right)^{\scriptscriptstyle T},\qquad|\beta\rangle=\left(\beta_{1},...,\beta_{N}\right)^{\scriptscriptstyle T},$ (3.3)
while $N$-rows $\langle a|$ and $\langle b|$
$\langle a|=\left(a_{1},...,a_{N}\right),\qquad\langle b|=\left(b_{1},...,b_{N}\right)$ (3.4)
depend on the coordinates and, in turn, determine the coordinate dependence of the $N\times N$ matrices $\mathsf{A}$ and $\mathsf{B}$.
The recipe for soliton solutions consists of two parts. The first step is to ‘construct’ functions $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ of the matrices and vectors introduced above. As was mentioned in the Introduction, we use the already known results for the ALH. So, we follow section 2.2 of the paper [8], and present functions (2.6) as
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle a|\mathsf{R}^{-1}\mathsf{F}|\beta\rangle,$ (3.5)
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\langle b|\mathsf{L}^{-1}\mathsf{G}|\alpha\rangle,$ (3.6)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|\alpha\rangle,$ (3.7)
$\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1+\langle b|\mathsf{L}^{-1}\mathsf{G}\mathsf{A}|\beta\rangle$ (3.8)
where
$\mathsf{F}=\left(1+\mathsf{B}\mathsf{A}\right)^{-1},\qquad\mathsf{G}=\left(1+\mathsf{A}\mathsf{B}\right)^{-1}.$ (3.9)
The next step, is the dependence of $\mathsf{A}$, $\mathsf{B}$ etc on the variables of the hierarchy $z_{1}$, $z_{2}$, … (and, in particular, on $x$ and $t$). Again, we do not invent anything new and use the ‘classical’ prescription. The case is that usually, in soliton solutions, the $z_{k}$-dependence appears through various exponential functions like $\exp[-i\phi(\boldsymbol{z})]$ with $\phi(\boldsymbol{z})=\sum_{k=1}^{\infty}c^{k}z_{k}$ (the simplest non-trivial series). The importance of such functions stems from the identity
$\mathbb{E}_{\xi}{\phi}-\phi=i\sum_{k=1}^{\infty}\frac{(c\,\xi)^{k}}{k}=-i\ln(1-c\xi)$ (3.10)
and, as a result, from the fact that $\exp[-i\phi(\boldsymbol{z})]$ is an eigenfunction of the shift operator $\mathbb{E}_{\xi}$:
$\mathbb{E}_{\xi}{e^{-i\phi(\boldsymbol{z})}}=e^{-i\phi(\boldsymbol{z})}/(1-c\xi).$ (3.11)
Considering our problem, we, as in [8], define
$\mathbb{E}_{\xi}\langle a|=\langle a|\mathsf{J}_{\xi}^{-1},\qquad\mathbb{E}_{\xi}\langle b|=\langle b|\mathsf{K}_{\xi}$ (3.12)
with
$\mathsf{J}_{\xi}=1-\xi\mathsf{R}^{-1},\qquad\mathsf{K}_{\xi}=1-\xi\mathsf{L}^{-1}$ (3.13)
(a vector version of (3.11)) which clearly implies
$\mathbb{E}_{\xi}\mathsf{A}=\mathsf{A}\mathsf{J}_{\xi}^{-1},\qquad\mathbb{E}_{\xi}\mathsf{B}=\mathsf{B}\mathsf{K}_{\xi}$ (3.14)
and then prove in Appendix B that functions (3.5)–(3.8) satisfy (2.7) and (2.8). Moreover, these functions are solutions of all differential equations of the ALH (equations (2.13)–(2.16) included), provided the variables of the hierarchy, $z_{1},z_{2},$..., are introduced in accordance with the definition (3.12). Returning to our problem, this means that the $(x,t)$-dependency is governed by
$i\mathsf{A}_{x}=\mathsf{A}\mathsf{R}^{-1},\qquad i\mathsf{A}_{t}=\mathsf{A}\mathsf{R}^{-2}$ (3.15)
and
$i\mathsf{B}_{x}=-\mathsf{B}\mathsf{L}^{-1},\qquad i\mathsf{B}_{t}=-\mathsf{B}\mathsf{L}^{-2},$ (3.16)
with similar equations for $\langle a|$ and $\langle b|$. Summarizing, we can formulate the following result.
Proposition 3.1
Vectors $\boldsymbol{q}$ and $\boldsymbol{r}$ defined in (2.17) and (3.5)–(3.8) with
$\langle a(x,t)|=\langle a_{0}|\mathsf{E}_{A}(x,t),\quad\mathsf{A}(x,t)=\mathsf{A}_{0}\mathsf{E}_{A}(x,t),$ (3.17)
$\langle b(x,t)|=\langle b_{0}|\mathsf{E}_{B}(x,t),\quad\mathsf{B}(x,t)=\mathsf{B}_{0}\mathsf{E}_{B}(x,t),$ (3.18)
where $\langle a_{0}|=\left(a_{01},...,a_{0N}\right)$ and $\langle b_{0}|=\left(b_{01},...,b_{0N}\right)$ are arbitrary constant NN-rows,
$\mathsf{A}_{0}=\Biggl(\frac{\alpha_{j}a_{0k}}{L_{j}-R_{k}}\Biggr)_{j,k=1,...,N},\qquad\mathsf{B}_{0}=\Biggl(\frac{\beta_{j}b_{0k}}{R_{j}-L_{k}}\Biggr)_{j,k=1,...,N}$ (3.19)
and
$\mathsf{E}_{A}(x,t)=\exp\left(-ix\mathsf{R}^{-1}-it\mathsf{R}^{-2}\right),$ (3.20)
$\mathsf{E}_{B}(x,t)=\exp\left(ix\mathsf{L}^{-1}+it\mathsf{L}^{-2}\right)$ (3.21)
satisfy equations (2.21)–(2.24) and constraint (2.20).
Thus, we have derived soliton solutions for the ‘two-field’ version of our problem, consisting of linear Schrödinger equations under the bilinear constraint (2.20).
The last step is to pass from solutions described in Proposition 3.1 to solution of our problem. We start with imposing some restrictions on the constants involved to ensure the relations
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}{\!}^{*},\qquad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}{\!}^{*}.$ (3.22)
It turns out that this can be achieved by taking
$R_{j}=L_{j}^{*},\quad\beta_{j}a_{0j}=(\alpha_{j}b_{0j})^{*},\qquad j=1,...,N.$ (3.23)
This, at first, implies $\mathsf{E}_{B}=\mathsf{E}_{A}^{*}$. Then, after rewriting (3.19) as $\mathsf{A}_{0}=\mathsf{D}_{\alpha}\mathsf{C}\mathsf{D}_{a}$ and $\mathsf{B}_{0}=\mathsf{D}_{\beta}\mathsf{C}^{*}\mathsf{D}_{b}$, where
$\mathsf{C}=\Biggl(\frac{1}{L_{j}-L_{k}^{*}}\Biggr)_{j,k=1,...,N}$ (3.24)
and $\mathsf{D}_{\alpha}$, $\mathsf{D}_{\beta}$, $\mathsf{D}_{a}$, $\mathsf{D}_{b}$ are diagonal matrices with elements $\alpha_{j}$, $\beta_{j}$, $a_{0j}$, $b_{0j}$ (j=1,,Nj=1,...,N) correspondingly, one can present $\mathsf{A}$ and $\mathsf{B}$ as
$\mathsf{A}=\mathsf{D}_{\alpha}\mathsf{Y}\mathsf{D}_{\beta}^{-1},\quad\mathsf{B}=\mathsf{D}_{\beta}\mathsf{Y}^{*}\mathsf{D}_{\alpha}^{-1}$ (3.25)
where
$\mathsf{Y}=\mathsf{C}\mathsf{D}_{\beta}\mathsf{D}_{a}\mathsf{E}_{A}.$ (3.26)
which leads to
$\mathsf{F}=\mathsf{D}_{\beta}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1}\mathsf{D}_{\beta}^{-1},\qquad\mathsf{G}=\mathsf{D}_{\alpha}\left(\mathsf{1}+\mathsf{Y}\mathsf{Y}^{*}\right)^{-1}\mathsf{D}_{\alpha}^{-1}.$ (3.27)
After some simple calculations, one can rewrite $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ as
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\gamma|\mathsf{\Omega}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}=1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle$ (3.28)
where
$\mathsf{\Omega}=\mathsf{Y}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},$ (3.29)
the constant row $\langle\gamma|$ is given by
$\langle\gamma|=\langle 1|\left(\mathsf{C}\mathsf{L}^{*}\right)^{-1}$ (3.30)
and
$\langle 1|=\left(1,...,1\right),\quad|1\rangle=\left(1,...,1\right)^{T}.$ (3.31)
In a similar way, the ansatz for $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$, given by (3.6) and (3.8), rewritten in terms of $\mathsf{Y}$, $\mathsf{\Omega}$ and $\langle\gamma|$ leads to
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\langle\gamma^{*}|\mathsf{\Omega}^{*}|1\rangle,\quad\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=1+\langle\gamma^{*}|\mathsf{\Omega}^{*}\mathsf{Y}|1\rangle$ (3.32)
which demonstrates that $\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}^{\!*}$ and $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}^{\!*}$.
The dependence on $t$ and $x$ is described by the matrix $\mathsf{E}$,
$\mathsf{E}=\mathsf{D}_{\beta}\mathsf{D}_{a}\mathsf{E}_{A}$ (3.33)
which can be written as
$\mathsf{E}=\mathop{\mbox{diag}}\left(\;e^{f_{k}+i\varphi_{k}}\;\right)_{k=1,...,N}$ (3.34)
with
$\begin{array}[]{lcl}f_{k}(t,x)&=&\nu_{k}x+2\mu_{k}\nu_{k}t+f_{0k},\\[5.69054pt] \varphi_{k}(t,x)&=&-\mu_{k}x+(\nu_{k}^{2}-\mu_{k}^{2})t+\varphi_{0k}\end{array}$ (3.35)
where $\mu_{k}$ and $\nu_{k}$ are defined by
$\mu_{k}+i\nu_{k}=L_{k}\left/|L_{k}|^{2}\right$. (3.36)
and $f_{0k}$, $\varphi_{0k}$ are arbitrary constants.
Now, we have all necessary to formulate the main result of this paper.
Proposition 3.2
Vectors $\boldsymbol{\psi}$ defined by
$\boldsymbol{\psi}=\mathsf{U}\left(\begin{array}[]{c}\langle\gamma|\mathsf{\Omega}|1\rangle,\quad\\ 1+\langle\gamma|\mathsf{\Omega}\mathsf{Y}^{*}|1\rangle\end{array}\right)$ (3.37)
where $\mathsf{U}$ is an arbitrary constant unitary matrix,
$\mathsf{\Omega}=\mathsf{Y}\left(\mathsf{1}+\mathsf{Y}^{*}\mathsf{Y}\right)^{-1},\quad\mathsf{Y}=\mathsf{C}\mathsf{E},$ (3.38)
with $\langle\gamma|$, $\mathsf{C}$ and $\mathsf{E}$ defined in (3.30), (3.24) and (3.34), solve the Euler–Lagrange equations (1.3). Elements of the matrix $\mathsf{U}$ together with the $4N$ constants $\mathrm{Re}\,L_{k}$, $\mathrm{Im}\,L_{k}$, $f_{0k}$ and $\varphi_{0k}$ are arbitrary parameters that determine the properties of the $N$-soliton solution.
To get some insight into the structure of obtained solutions, let us consider the one-soliton solution.
In this case the matrices $\mathsf{L}$, $\mathsf{C}$ and the rows $\langle\gamma|$ become just scalars,
$\mathsf{L}\to L_{1},\quad\mathsf{C}\to\frac{1}{2i|L_{1}|\sin\theta},\quad\langle\gamma|\to e^{2i\theta}-1$ (3.39)
where $\theta=\arg L_{1}$, or
$\theta=\tan^{-1}(\nu_{1}/\mu_{1}).$ (3.40)
Modifying slightly the functions $f_{1}$ and $\varphi_{1}$ that appear in (3.34), i.e. introducing the new ones, $f$ and $\varphi$, defined by
$f(t,x)=f_{1}(t,x)+\ln\frac{\mu_{1}^{2}+\nu_{1}^{2}}{2|\nu_{1}|},$ (3.41)
$\varphi(t,x)=\varphi_{1}(t,x)-\frac{\pi}{2}\mathop{\mbox{sign}}\nu_{1}$ (3.42)
one can present $\mathsf{Y}$ as $\mathsf{\Omega}$ as
$\mathsf{Y}=\exp(f+i\varphi),\quad\mathsf{\Omega}=\frac{1}{2}\exp(i\varphi)\mathop{\mbox{sech}}f.$ (3.43)
Choosing $\mathsf{U}=\mbox{diag}\left(e^{-i\theta},e^{-i\theta}\right)$ one can obtain the following expressions for the components of the vector $\boldsymbol{\psi}=\left(\psi_{1},\psi_{2}\right)^{\scriptscriptstyle T}$:
$\psi_{1}=i\sin\theta\,\exp(i\varphi)\mathop{\mbox{sech}}{f},$ (3.44)
$\psi_{2}=\cos\theta+i\sin\theta\,\tanh f.$ (3.45)
with
$f(t,x)=\nu_{1}x+2\mu_{1}\nu_{1}t+f_{0},$ (3.46)
$\varphi(t,x)=-\mu_{1}x+(\nu_{1}^{2}-\mu_{1}^{2})t+\varphi_{0}.$ (3.47)
It is easy to see that $\psi_{1}$ and $\psi_{2}$ are the NLSE-solitons of different type: $\psi_{1}$ is a so-called bright soliton, vanishing as $x\to\pm\infty$ (with $t$ being fixed), while $\psi_{2}$ is a dark soliton ($\lim_{|x|\to\infty}\psi_{2}=e^{\pm i\theta}$). However, this does not mean that the same is true for any solution (even for the one-soliton one), because for an arbitrary $\mathsf{U}$, all components of the general $N$-soliton solution are mixtures of dark and bright solitons.
Another useful information that can be obtained form (3.44)–(3.47) is the fact that all physically important characteristics of the soliton, its amplitude ($\sin\theta$), velocity ($2\mu_{1}$) and the scale ($1/\nu_{1}$), are determined by the choice of $L_{1}$. The same is true in the multi-soliton case. Although a $N$-soliton solution of a nonlinear equation is surely not a sum of $N$ solitons (only in the asymptotic regions, $t\to\pm\infty$ it can be viewed as such), it is possible to describe its structure qualitatively in terms of single solitons whose amplitude, velocity and scale depend on on the elements of the matrix $\mathsf{L}$, $L_{k}$, while other parameters, $\alpha_{k}$, $\beta_{k}$, $a_{0k}$ and $b_{0k}$, which were ‘absorbed’ into the matrix $\mathsf{E}$ (3.34), i.e. replaced by $f_{0k}$ and $\varphi_{0k}$, determine their relative position and phasing. The role of the matrix $\mathsf{U}$ is mostly ‘mixing’ of different solitons.
4 Vector NLSE with gradient nonlinearity.
In this section we consider an example of application of the results, obtained in this paper for a rather abstract system, to a more physical problem.
Let us return to the system (2.21), (2.22) and study the behavior of the function $s$ defined as
$s=\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{x}=-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}$ (4.1)
(the last equality stems from the fact $\partial_{x}(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q})=0$). Calculating the derivative of $s$ using both of the equations in (4.1) and expressing $\boldsymbol{q}_{xx}$ and $\boldsymbol{r}^{\scriptscriptstyle T}_{xx}$ from (2.21) and (2.22), one can easily obtain
$s_{x}$ $=$ $-i\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}-\lambda$ (4.2)
$=$ $-i\boldsymbol{r}^{\scriptscriptstyle T}_{t}\boldsymbol{q}-\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}+\lambda$ (4.3)
which, after summation of both expressions, leads to
$s_{x}=-(i/2)\,\partial_{t}(\boldsymbol{r}^{\scriptscriptstyle T}\boldsymbol{q})=0$ (4.4)
which means that $s=s(t)$.
In a similar way, one can calculate the time derivative of $s$ which leads to
$is_{t}=\partial_{x}\left(2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}-\lambda\right).$ (4.5)
This implies that
$\lambda=\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}$ (4.6)
where
$\lambda_{0}=-is^{\prime}(t)x+s_{1}(t)$ (4.7)
and $s_{1}$ is another functions not depending on $x$. Thus, equations (2.21) and (2.22) can be rewritten as
$i\boldsymbol{q}_{t}+\boldsymbol{q}_{xx}+(\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x})\boldsymbol{q}=0,$ (4.8)
$i\boldsymbol{r}^{\scriptscriptstyle T}_{t}+\boldsymbol{r}^{\scriptscriptstyle T}_{xx}+(\lambda_{0}+2\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x})\boldsymbol{r}^{\scriptscriptstyle T}=0$ (4.9)
In the case of the soliton solutions presented in the previous section, it can be shown, by calculating $\boldsymbol{r}^{\scriptscriptstyle T}_{x}\boldsymbol{q}_{x}$ using equations (2.18) and comparing the result with the expression for $\lambda$ in (2.23), that $\lambda_{0}=0$. Thus, as a byproduct of the calculations presented in this paper, we obtain the following result.
Proposition 4.1
NN-soliton solutions described in proposition 3.2 are, at the same time, solutions of the NLSE with gradient nonlinearity
$i\boldsymbol{\psi}_{t}+\boldsymbol{\psi}_{xx}+2\,\boldsymbol{\psi}^{\dagger}_{x}\boldsymbol{\psi}_{x}\,\boldsymbol{\psi}=0.$ (4.10)
5 Conclusion.
In this work we have established the relationship between the Schrödinger equation with the constraint and the ALH. As was mentioned in section 2, we considered only the ‘positive’ subhierarchy (2.4). As to the ‘negative’ subhierarchy, it can be shown that calculations similar to ones presented above lead to the set of solutions similar to the solutions described in the proposition 3.2.
A more interesting question is whether we can tackle with the approach of this paper the case on quadratic restrictions other than (1.2), for example, ones given by
$\boldsymbol{\psi}^{\dagger}\sigma_{3}\boldsymbol{\psi}=1$ (5.1)
where $\sigma_{3}=\mbox{diag}\left(1,-1\right)$? The answer, which we present here without derivation, is ‘yes’. However, to do this one should start not with the bright solitons of the ALH, as in section 3, but with the dark ones. In some sense the signature of the matrix describing the applied constraints play the role of the sign in front of the nonlinear term in the NLSE: it determines which kind solitons (bright or dark) exists in the system.
Finally we would like to add a short comment on the integrability of the system (1.3) which was not discussed in the paper. We cannot at present prove its integrability, for example, by developing the inverse scattering transform. However, we now know that equations (1.3) possess $N$-soliton solutions. That means that the model (1.1) passes the so-called $N$-soliton test for integrability, which, though not proved rigorously, has a long history of successful applications and was even used as a tool for finding new integrable models, as, for example, in the comprehensive study by Hietarinta [9, 10, 11, 12]. Also, we now know that equations (1.3) are a consequence of equations of the integrable ALH. These two facts are a strong indication that the model (1.1) is integrable. Nevertheless, we think that the work in this direction should be continued, and one of the first problems to solve is to find the conservation laws of the model, which may be a topic of the following study.
Acknowledgments.
We would like to thank the Referees for careful reading of the manuscript and their valuable comments and suggestions, which helped to improve this paper.
Appendix A. A proof of (2.26).
To prove the fact that functions $u$ and $v$ satisfy the NLSE we need some consequences of the formulae presented in section 2.
First, it follows from equations (2.13) and (2.14) together with the restriction (2.10), that functions $u$ and $v$ can be written as
$u=i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}-i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x},\qquad v=i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}-i\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}$ (A.1)
and that
$uv=\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}.$ (A.2)
Differentiating (A.1) with respect to $t$ and expressing the $t$-derivatives from (2.15) and (2.16) leads to
$\begin{array}[]{lcl}u_{t}&=&-u_{xx}+\left(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\right)u_{x}\\ &&+\left(\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{xx}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{xx}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}\right)u.\end{array}$ (A.3)
Another consequence of (2.13) and (2.14) is that
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}=0,\qquad\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}=0$ (A.4)
which, in particular, implies
$\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{xx}+\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{xx}=-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}_{x}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}_{x}\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}_{x}.$ (A.5)
Thus, the first braces in the right-hand side of equation (A.3) disappear while the factor in front of $u$ becomes just 2uv-2uv. Equation (A.3) now reads
$u_{t}=-u_{xx}-2u^{2}v$ (A.6)
which is nothing but the first equation from (2.26).
The second equation from (2.26) can be demonstrated in the similar way.
Appendix B. Validation of the ansatz.
Here we demonstrate that ansatz (3.5)–(3.8) leads to the solutions of the (2.7) and (2.8)
As follows from (3.5) and (3.12),
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|\beta\rangle-\langle{a}|\mathsf{R}^{-1}\mathsf{F}|\beta\rangle$ (B.1)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})|\beta\rangle-\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}\mathsf{J}_{\xi}\mathsf{F}|\beta\rangle$
which can be rewritten as
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}=\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}|\beta\rangle$ (B.2)
where $\mathsf{X}$ is defined by
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{X}\mathsf{F}=\mathbb{E}_{\xi}{\mathsf{F}}-\mathsf{J}_{\xi}\mathsf{F}.$ (B.3)
Using the definition of $\mathsf{F}$ (3.9), (3.14) and then (3.1) one can obtain
$\mathsf{X}$ $=$ $1-\mathsf{J}_{\xi}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}$ (B.4)
$=$ $\xi\mathsf{R}^{-1}+\mathsf{B}\mathsf{A}-(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}$
$=$ $\xi\mathsf{R}^{-1}+\xi\mathsf{B}\mathsf{L}^{-1}\mathsf{A}$
$=$ $\xi\mathsf{R}^{-1}\mathsf{F}^{-1}+\xi\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}.$
Substituting this expression into (B.2) and using (3.8) together with the identity $\mathsf{A}\mathsf{F}=\mathsf{G}\mathsf{A}$, one arrives at
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{q}}$ $=$ $\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle+\xi\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1}\mathsf{A}\mathsf{F}|\beta\rangle$ (B.5)
$=$ $\xi\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{r}}\mathbb{E}_{\xi}{{u}}$
which is nothing but the first equation from (2.7).
In a similar way one can prove that $\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ satisfies the first equation from (2.8).
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})|\alpha\rangle-\langle a|\mathsf{R}^{-1}\mathsf{F}\mathsf{B}|\alpha\rangle$ (B.6)
$=$ $\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}|\alpha\rangle$
where
$(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{Y}\mathsf{G}=(\mathbb{E}_{\xi}{\mathsf{F}})(\mathbb{E}_{\xi}{\mathsf{B}})-\mathsf{J}_{\xi}\mathsf{F}\mathsf{B}.$ (B.7)
Calculating $\mathsf{Y}$,
$\mathsf{Y}$ $=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}+(\mathbb{E}_{\xi}{\mathsf{B}})\mathsf{A}\mathsf{B}-(\mathbb{E}_{\xi}{\mathsf{B}})(\mathbb{E}_{\xi}{\mathsf{A}})\mathsf{J}_{\xi}\mathsf{B}$ (B.8)
$=$ $\mathbb{E}_{\xi}{\mathsf{B}}-\mathsf{J}_{\xi}\mathsf{B}$
$=$ $\mathsf{B}\mathsf{K}_{\xi}-\mathsf{J}_{\xi}\mathsf{B}$
$=$ $\xi\mathsf{R}^{-1}\mathsf{B}-\xi\mathsf{B}\mathsf{L}^{-1}$
$=$ $-\xi\mathsf{R}^{-1}|\beta\rangle\langle b|\mathsf{L}^{-1},$
and substituting it in (B.6) one can obtain
$\mathbb{E}_{\xi}{\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}}-\stackrel{{\scriptstyle\scriptscriptstyle 2}}{{q}}$ $=$ $-\xi\langle b|\mathsf{L}^{-1}\mathsf{G}|\alpha\rangle\langle\mathbb{E}_{\xi}{{a}}|\mathsf{R}^{-1}(\mathbb{E}_{\xi}{\mathsf{F}})\mathsf{R}^{-1}|\beta\rangle$ (B.9)
$=$ $-\xi\stackrel{{\scriptstyle\scriptscriptstyle 1}}{{r}}\mathbb{E}_{\xi}{{u}},$
which concludes the proof.
The rest of the equations (2.7) and (2.8) can be tackled in a similar way.